# Exploring the range of impacts of helium in the spectra of double detonation models for Type Ia supernovae

### v1

@@ 5 Conclusions
We have carried out a NLTE radiative transfer simulation including treatment for non-thermal electrons for a realisation of the double detonation scenario with a modest mass of He that is ejected from the outer helium layer with relatively low velocities. We find our model predicts a strong and persistent He i 10830 Å spectral feature that is is strongly blended with the spectral contribution of Mg ii 10927 Å. This Mg ii feature shows an extended wing in the spectra of some normal SNe Ia that has previously been attributed to C i 10693 Å (Hsiao et al., 2013; Marion et al., 2015). However, the He i 10830 Å feature predicted by our simulation has its strongest absorption to the blue wing of this Mg ii 10927 Å feature and our model does not predict a C i 10693 Å feature forms. We therefore suggest the extended wing may instead be a spectral signature of He, although we note the He feature predicted by our particular model is too strong and persistent to be consistent with normal SNe Ia.
Despite the clear He i 10830 Å feature feature predicted by our simulation our model does not predict any clear optical He spectral features or a clear 2 m feature. Therefore, consistent with what C23 found for their double detonation simulation, our simulation demonstrates that the non-detection of clear optical He features can not be used to rule out the presence of a He i 10830 Å spectral feature.
Compared to the C23 double-detonation model, which predicted a distinct He i 10830 Å feature blueward of Mg ii 10927 Å, the model investigated in this study has a substantially higher mass of ejected He, unburnt He present at lower velocities and a significantly lower mass of in the He-rich outer region. Therefore, together these simulations demonstrate that He i 10830 Å can exhibit different spectral signatures depending on the exact realisation of the double detonation scenario. In particular the formation of the He i 10830 Å feature, and therefore the exact nature of the He spectral signature predicted, appears to depend strongly on the distribution of unburnt He in the ejecta which can vary significantly between double detonation simulations, and with line-of-sight in multi-dimensional models.
@@
Future work should focus on full NLTE multi-dimensional radiative transfer simulations of double detonation models that will allow the viewing angle variation of predicted He spectral signatures to be explored. Such simulations will also be key to determining whether realisations of the double detonation scenario exist that predict He spectral signatures consistent with the NIR spectra of normal SNe Ia.
@@ Acknowledgements
FPC and SAS, acknowledge funding from STFC grant ST/X00094X/1. This work used the DiRAC Memory Intensive service (Cosma8) at Durham University, managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The DiRAC service at Durham was funded by BEIS, UKRI and STFC capital funding, Durham University and STFC operations grants. Access to DiRAC resources was granted through a Director’s Discretionary Time allocation in 2023/24. This work also used the DiRAC Data Intensive service (CSD3) at the University of Cambridge, managed by the University of Cambridge University Information Services on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The DiRAC component of CSD3 at Cambridge was funded by BEIS, UKRI and STFC capital funding and STFC operations grants. DiRAC is part of the UKRI Digital Research Infrastructure. The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC). CEC acknowledges funding by the European Union (ERC, HEAVYMETAL, 101071865). LJS acknowledges support by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (ERC Advanced Grant KILONOVA No. 885281). LJS acknowledges support by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Project-ID 279384907 - SFB 1245 and MA 4248/3-1. JMP acknowledges the support of the Department for Economy (DfE). The work of FKR is supported by the EuropeanUnion (ERC, ExCEED, 101096243) and by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - project number 537700965. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. This work was supported in part by the European Union (ChETEC-INFRA, project no. 101008324), NSF/IReNA and the Klaus Tschira Foundation. NumPy and SciPy (Oliphant, 2007), Matplotlib (Hunter, 2007) and artistools222https://github.com/artis-mcrt/artistools/ (Shingles et al., 2023) were used for data processing and plotting.

### v2

@@ 4 Conclusions
We have carried out a NLTE radiative transfer simulation including treatment for non-thermal electrons for a realisation of the double detonation scenario with a modest mass of He (0.04 ) that is ejected from the outer helium layer with relatively low velocities (). The model is representative of the negative polar direction in the 3D Gronow et al. (2021) model on which it is based (where the He detonation is ignited in the positive polar direction). We find our model predicts a strong and persistent He i 10830 Å spectral feature that is strongly blended with the spectral contribution of Mg ii 10927 Å. This Mg ii feature shows an extended wing in the spectra of some normal SNe Ia that has previously been attributed to C i 10693 Å (Hsiao et al., 2013; Marion et al., 2015). However, the He i 10830 Å feature predicted by our simulation has its strongest absorption to the blue wing of this Mg ii 10927 Å feature and our model does not predict a C i 10693 Å feature forms. We therefore suggest the extended wing may instead be a spectral signature of He, although we note the He feature predicted by our particular model is too strong and persistent to be consistent with normal SNe Ia.
Despite the strong He i 10830 Å feature feature predicted by our simulation no clear optical He spectral features appear. Therefore, consistent with what C23 found for their double detonation simulation, our simulation demonstrates that the non-detection of optical He features can not be used to rule out the He i 10830 Å spectral feature. Our simulation also predicted no clear He i 2 m feature forms, demonstrating the absence of He i 2 m also does not rule out He i 10830 Å (the C23 simulation did not have sufficient signal to noise to test the formation of the He i 2 m feature). The He mass of our model is likely towards the high end of what can be ejected while still producing reasonable agreement with the optical spectra of normal and transitional SNe Ia (Kromer et al., 2010) which suggests observed He i 10830 Å spectral signatures present in the spectra of normal or transitional SNe Ia are not necessarily accompanied by additional He signatures.
Relative to our simulation, the double detonation model presented by C23 (which also included a full treatment of NLTE and non-thermal processes) has a significantly lower He ejecta mass (0.018 ), present at higher velocities () and is representative of the equatorial direction in the 3D Gronow et al. (2020) model on which it is based. The contribution of He i 10830 Å is weaker and present at higher velocities for the C23 simulation, forming a distinct feature blueward of Mg ii 10927 Å which is no longer clearly present by peak. This feature is consistent with a feature present in the pre-peak spectra of observed transitional SNe Ia (Hsiao et al., 2015; Wyatt et al., 2021; Pearson et al., 2024). Together these simulations demonstrate that the He signatures predicted by double detonation models can show substantial variation depending on the exact realisation of the ejecta structure, and in particular the distribution of unburnt He, which is both model and line-of-sight dependent. This implies significant diversity is expected in the He i 10830 Å spectral signature across the observed SN Ia population. This should be tested with observations of the pre-peak NIR spectroscopic evolution for a larger sample of SNe Ia.
@@
Future work should also focus on full NLTE multi-dimensional radiative transfer simulations of double detonation models that will allow the viewing angle variation of predicted He spectral signatures to be explored. In particular, since both our simulation and the C23 simulation predict a He i 10830 Å spectral signature too strong compared to normal SNe Ia, double detonation models with lower masses of ejected He should be prioritised. Such simulations will be key to determining whether realisations of the double detonation scenario exist that predict He spectral signatures consistent with the NIR spectra of normal SNe Ia and will also allow us to probe the lower limit of He that can be ejected while still producing an observable signature.
@@ Acknowledgements
FPC and SAS, acknowledge funding from STFC grant ST/X00094X/1. This work used the DiRAC Memory Intensive service (Cosma8) at Durham University, managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The DiRAC service at Durham was funded by BEIS, UKRI and STFC capital funding, Durham University and STFC operations grants. Access to DiRAC resources was granted through a Director’s Discretionary Time allocation in 2023/24. This work also used the DiRAC Data Intensive service (CSD3) at the University of Cambridge, managed by the University of Cambridge University Information Services on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The DiRAC component of CSD3 at Cambridge was funded by BEIS, UKRI and STFC capital funding and STFC operations grants. DiRAC is part of the UKRI Digital Research Infrastructure. The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC). CEC acknowledges funding by the European Union (ERC, HEAVYMETAL, 101071865). LJS acknowledges support by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (ERC Advanced Grant KILONOVA No. 885281). LJS acknowledges support by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Project-ID 279384907 - SFB 1245 and MA 4248/3-1. JMP acknowledges the support of the Department for Economy (DfE). The work of FKR is supported by the EuropeanUnion (ERC, ExCEED, 101096243) and by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - project number 537700965. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. This work was supported in part by the European Union (ChETEC-INFRA, project no. 101008324), NSF/IReNA and the Klaus Tschira Foundation. NumPy and SciPy (Oliphant, 2007), Matplotlib (Hunter, 2007) and artistools222https://github.com/artis-mcrt/artistools/ (Shingles et al., 2025) were used for data processing and plotting.